# Threshold Graphs are Globally Synchronizing

### v4

@@ 3.1. Equilibrium and Local Stability
$\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\boldsymbol{\theta}_{j}-\boldsymbol{\theta}_{i})=0,\quad\forall i\in[n].$ (5)
An equilibrium point is considered stable if small perturbations around it remain small over time [Kha02]. The linearization of a nonlinear system provides a necessary condition for local stability: an equilibrium point is stable if the Jacobian matrix is negative semi-definite.
Proposition 3.2 (Characterization of stable equilibrium).
An equilibrium point $\boldsymbol{\theta}$ is stable if the Jacobian matrix $\boldsymbol{\mathbf{J}}$, defined as:
@@
3.2. Stationary Point and Local Optimality
Definition 3.3 (Stationary point of the energy function (4)).
We call $\theta\in\mathbb{R}^{n}$ a stationary point of the energy function (4) if it satisfies $\nabla E(\theta)=0$, namely
$\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\boldsymbol{\theta}_{j}-\boldsymbol{\theta}_{i})=0,\quad\forall i\in[n].$ (6)
Obviously, from Proposition 3.1 and Definition 3.3, a configuration $\boldsymbol{\theta}\in\mathbb{R}^{n}$ is a stationary point of the energy function (4) if and only if it is an equilibrium of the Kuramoto model (2).
@@ 3.2. Stationary Point and Local Optimality
A configuration $\theta\in\mathbb{R}^{n}$ is called a local minimum of the energy function $E$ if there exists $\varepsilon>0$ such that
$E(\theta)\leq E(\theta^{\prime}),\quad\forall\theta^{\prime}\in\mathbb{R}^{n}\ \text{with}\ \|\theta^{\prime}-\theta\|<\varepsilon.$
Proposition 3.5 (First- and second-order necessary conditions).
If a configuration $\theta$ is a local minimum of $E$, then
  1. (1)
    $\nabla E(\theta)=0$ (first-order condition), and
  2. (2)
    the Hessian $\nabla^{2}E(\theta)$ is positive semidefinite.
A point satisfying the above two conditions is also called a second-order stationary point (SOSP) of (4).
3.3. Correspondence between Stability and Optimality
In this section, we briefly outline the relationship between the equilibrium stability of the dynamical system (2) and the local optimality of the energy function (4). Understanding this correspondence provides valuable intuition. In particular, in this work, we often exploits the fact that a small perturbation causes the system to leave that state (therefore this state in unstable), in order to show a certain state is not a second-order stationary point.
At the first-order level, the two notions coincide. Indeed, a configuration $\boldsymbol{\theta}$ is a first-order stationary point of the energy function (4) if and only if it is an equilibrium of the dynamical system (2) on the graph $G$ with adjacency matrix $A$, since both are characterized by
$\sum_{j=1}^{n}A_{ij}\sin(\theta_{j}-\theta_{i})=0\qquad\text{for all }i\in V.$
One can further relate the stability of an equilibrium to the second-order necessary condition of the energy landscape. It is precisely because of the real analycity of the energy function (4), the local stability of an equilibrium of (2) is equivalent to the local minimality of the energy function (4). The relation mentioned here is summarized in Figure 4. Such an equivalence was established in the celebrated work [AK06].
@@ 4. Vector Geometry of Equilibria
In this section, we recast the algebraic first- and second-order stationary conditions in Section 3 into geometric ones. This geometric viewpoint replaces the nonlinear equations (6) and the Hessian constraints by combinatorial relations among the vectors associated with the nodes. They form the starting point of our analysis. Additionally, stronger geometric constraints will be derived in later sections and combined with the results here to rule out all non-synchronous second-order stationary points of the energy function $E_{G}(\boldsymbol{\theta})$ where $G$ is any connected threshold graph.
@@ 4.1. Vector-labeled Graph
The basic object underlying this geometric perspective is the vector-labeled graph corresponding to angle configuration of $\boldsymbol{\theta}$. Defining such object allows us reasoning about whether a given state $\boldsymbol{\theta}$ is an equilibrium, and whether an equilibrium is stable, i.e., can we move the angles so as to decrease the energy.
Given a state $\boldsymbol{\theta}\in\mathbb{R}^{n}$ of Kuramoto model on graph $G=(V,E)$. Associate each node $i\in V$ its corresponding unit vector
@@
This is called a phasor, following the standard convention in the Kuramoto model. The object $\left(G,\{\boldsymbol{v}_{i}\}_{i\in V}\right)$, called vector-labeled graph encoding $\boldsymbol{\theta}$ combinatorially and geometrically.
@@ 4.2. Alignment and Equilibrium
Given a vector-labeled graph $(G,\{\boldsymbol{v}_{i}\})$, we now formulate the equilibrium condition of (2) in geometric terms.
@@
where $v_{i}:=(\cos\theta_{i},\sin\theta_{i})$.
@@
Let us firstly derive an equivalent form of (7). Rewrite all vectors $\boldsymbol{v}_{i}$ in (7) in terms of their Cartesian components $(\cos\theta_{i},\sin\theta_{i})$, yielding
@@
This is equivalent to verifying the equality between two complex numbers:
@@
which can be compactly rewritten as
$\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}}.$ (9)
using the Euler formula $e^{I\theta_{i}}=\cos\theta_{i}+I\sin\theta_{i}$. Now that we have derived the equivalence between conditions (7) and (9), our goal now reduces to proving the equivalence between (9) and (4), namely
$\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}}\Longleftrightarrow\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})=0.$
This indeed holds.
To prove the implication from the right-hand side to the left-hand side, we construct such a $\mu_{i}$ that satisfies the left-hand side. Dividing both sides of (9) by $e^{I\theta_{i}}$ (noting that $e^{I\theta_{i}}\neq 0$), we obtain
$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}=\mu_{i}.$
Such a $\mu_{i}$ exists and is in fact given by
$\mu_{i}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i}),$
at an equilibrium point $\boldsymbol{\theta}$, as we will show below. Expanding the fraction, we obtain
$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}$ $=\sum_{j\in N(i)}e^{I(\theta_{j}-\theta_{i})}$ (10)
$=\sum_{j\in N(i)}(\cos(\theta_{j}-\theta_{i})+I\sin(\theta_{j}-\theta_{i})).$
Since $\boldsymbol{\theta}$ is an equilibrium, the imaginary part must vanish, leading to
$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})\in\mathbb{R}.$
Thus, we conclude
$\sum_{j\in N(i)}e^{I\theta_{j}}=\sum_{j\in N(i)}{\cos(\theta_{j}-\theta_{i})}\cdot e^{I\theta_{i}}.$
Conversely, to prove the left-hand side implies the right-hand side, we show that if there exists a constant $\mu_{i}$ such that
$\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}},$
then $\boldsymbol{\theta}$ is an equilibrium point. This is equivalent to showing that
@@
Indeed, we compute
$I\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})+\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})$ (11)
@@
Thus, the proof is complete. ∎
@@
Figure 6. Geometric intuition for the first- and second-order stationary conditions. The solid arrow represents $\boldsymbol{v}_{i}$, and the dashed arrow represents $\sum_{j\in N(i)}\boldsymbol{v}_{j}$. For each node $i$, the four panels show all possible configurations satisfying the first-order condition. Among them, cases (A) and (B) violate the second-order condition because the aggregated phasor lies outside the feasible (gray) region corresponding to $\mu_{i}\geq 0$.
4.3. Angle and Stability
Given an equilibrium $\boldsymbol{\theta}$, if there exists a node $i\in V$ whose corresponding force satisfies $\mu_{i}<0$, then $\boldsymbol{\theta}$ cannot be a second-order stationary point. Indeed, a perturbation that moves $\boldsymbol{v}_{i}$ toward the direction of $\sum_{j\in N(i)}\boldsymbol{v}_{j}$ strictly decreases the energy.
Lemma 4.4.
If $\boldsymbol{\theta}$ is a second-order stationary point of (4), then either $\sum_{j\in N(i)}\boldsymbol{v}_{j}=\boldsymbol{0}$ or
$\angle\!\left(\sum_{j\in N(i)}\boldsymbol{v}_{j},\,\boldsymbol{v}_{i}\right)=0.$
Proof of Lemma 4.4.
The $i$-th diagonal entry of $\mathrm{Hess}\,E(\theta)$ can be written as
$\mathrm{Hess}\,E(\boldsymbol{\theta})_{ii}:=\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j})=\sum_{j=1}^{n}\boldsymbol{A}_{ij}e^{I(\theta_{j}-\theta_{i})}=\mu_{i}$
where the second equality follows from the equilibrium condition (4.1), after rewriting the unit vectors in complex form. If some $\mu_{i}<0$, then $\mathrm{Hess}\,E(\theta)_{ii}<0$, contradicting the positive semidefiniteness of $\mathrm{Hess}\,E(\theta)$ at any second-order stationary point [Bha07] [HJ85]. ∎
Remark 4.5 (Mechanical insight into node at stable equilibrium).
The equilibrium condition in Lemma 4.4 for node $i\in V$ has a mechanical interpretation. The sum of the phasors representing its neighbors can be visualized as a set of unit-length rods with directions. We place one phasor’s starting point at the origin, then connect the remaining phasors head-to-tail. The head-to-tail connected rods can freely adjust their orientations, but the endpoint of the last phasor must remain on the right-end unrestricted sliding track, to ensure node $i$ being stable there.
Figure 7. With $\boldsymbol{v}_{i}$ fixed at angle 0, the endpoint of the aggregated phasor $\sum_{j}A_{ij}\boldsymbol{v}_{j}$ is constrained to the right-end sliding track. The rods may freely adjust their orientations, but the endpoint of the last rod must lie in the feasible sliding interval. The solid vectors show the unit-length phasors of the neighbors of node ii.
@@ 5.1. A Fundamental Geometric Fact
Let $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}\subset\mathbb{R}^{2}$. Suppose there exist a vector $\boldsymbol{q}\in\mathbb{R}^{2}$ and scalars $\mu_{a},\mu_{b}\in\mathbb{R}$ such that
@@
Case 1: $\mu_{a}=\mu_{b}=-1$. In this case, we have
$\boldsymbol{q}=\boldsymbol{v}_{a}+\boldsymbol{v}_{b}$
Case 2: $\mu_{a}\neq-1$ and $\mu_{b}\neq-1$. In this case, from (13), it must hold that
@@
This leads to two subcases: (1) If $\mu_{a}=\mu_{b}\neq-1$, then $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$; (2) If $\mu_{a}+1=-(\mu_{b}+1)$, then equivalently, we have $\mu_{a}+\mu_{b}=-2$. This implies $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$. ∎
@@
Figure 8. The feasible region for $(\mu_{a},\mu_{b})$ consists of: (1) the black line $\mu_{a}=\mu_{b}$, excluding the intersection point, corresponding to $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$; (2) the red line $\mu_{a}+\mu_{b}=-2$, excluding the intersection point, corresponding to the antipodal case $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$; and (3) their intersection point (1,1)(-1,-1), marked by a hollow dot, at which no additional constraint is imposed on the vectors $\boldsymbol{v}_{a}$ and $\boldsymbol{v}_{b}$.
@@
With $\mu_{a},\mu_{b}\geq 0$, Lemma 5.1 leaves only case (1), hence $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$. ∎
Definition 5.3 (Geometric twins).
At state $\boldsymbol{\theta}$ of Kuramoto model on $G$, nodes $i$ and $j$ are called geometric twins if their phasors $\boldsymbol{v}_{i}$ and $\boldsymbol{v}_{j}$ satisfy the compatibility relations of Lemma 5.1.

### v5

@@ 3.1. Equilibria and Local Stability
$\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i})=0,\quad\forall i\in[n].$ (3)
An equilibrium point is said to be stable if sufficiently small perturbations around it remain small for all future time [Kha02]. To assess the stability of a given equilibrium point, one typically studies the linearization of the system at that point, which is characterized by the Jacobian matrix.
Proposition 3.2 (Stable equilibrium of (1)).
An equilibrium point $\boldsymbol{\theta}\in\mathbb{R}^{n}$ of the Kuramoto model (1) is stable if the Jacobian matrix $\boldsymbol{\mathbf{J}}$, defined as:
@@
3.2. Stationary Points and Local Optimality
Definition 3.3 (First-order stationary points of (2)).
We call $\boldsymbol{\theta}\in\mathbb{R}^{n}$ a first-order stationary point of the energy function (2) if it satisfies $\nabla E(\boldsymbol{\theta})=0$, namely
$\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i})=0,\quad\forall i\in[n].$ (4)
Proposition 3.4 (Hessian $\boldsymbol{H}$ of (2)).
For the Kuramoto energy (2), the Hessian has entries
$\boldsymbol{H}_{ij}:=(\nabla^{2}E(\boldsymbol{\theta}))_{ij}=\begin{cases}\sum_{j=1}^{n}\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j}),&\text{if }i=j,\\[8.0pt] -\boldsymbol{A}_{ij}\cos(\theta_{i}-\theta_{j}),&\text{if }i\neq j,\end{cases}$
Definition 3.5 (Second-order stationary points of (2)).
A configuration $\boldsymbol{\theta}\in\mathbb{R}^{n}$ is called a second-order stationary point (SOSP) of (2) if $\nabla E(\boldsymbol{\theta})=0$, and the Hessian $\boldsymbol{H}$ is positive semidefinite.
3.3. Correspondence Between the Two Sets of Concepts
Proposition 3.6.
A configuration $\boldsymbol{\theta}\in\mathbb{R}^{n}$ is an equilibrium of the Kuramoto model (1) if and only if it is a first-order stationary point of the Kuramoto energy (2).
Proof.
This is clear, since both are characterized by
$\sum_{j=1}^{n}\boldsymbol{A}_{ij}\sin(\theta_{j}-\theta_{i})=0,\qquad\forall i\in[n].$
Moreover, since (2) is real analytic, an equilibrium of (1) is stable if and only if it is a local minimum of (2), modulo the global rotation symmetry. Such an equivalence was established in the celebrated work [AK06].
@@ 3.3. Correspondence Between the Two Sets of Concepts
A configuration $\boldsymbol{\theta}\in\mathbb{R}^{n}$ is called a local minimum of the energy function $E(\boldsymbol{\theta})$ if there exists $\varepsilon>0$ such that
$E(\boldsymbol{\theta})\leq E(\boldsymbol{\theta}^{\prime}),\quad\forall\boldsymbol{\theta}^{\prime}\in\mathbb{R}^{n}\ \text{with}\ \|\boldsymbol{\theta}^{\prime}-\boldsymbol{\theta}\|<\varepsilon.$
Therefore, the relations between stable equilibria, second-order stationary points and local minima can be summarized in Figure 4.
@@ 4. Vector Geometry of Equilibria
In this section, we recast the algebraic conditions introduced in Section 3 into geometric ones, namely combinatorial relations among vectors associated with the nodes. By doing so, the analysis of global synchronization on threshold graphs reduces to reasoning purely about geometric relations between the unit vectors corresponding to the angles in $\boldsymbol{\theta}$.
@@ 4.1. Vector-labeled Graph
Given a state $\boldsymbol{\theta}\in\mathbb{R}^{n}$ of the Kuramoto model on graph $G=(V,E)$. Associate each node $i\in V$ its corresponding unit vector
@@
This is called a phasor, following the terminology of Yoshiki Kuramoto. The object $\left(G,\{\boldsymbol{v}_{i}\}_{i\in V}\right)$, called vector-labeled graph encoding $\boldsymbol{\theta}$ combinatorially and geometrically.
@@ 4.2. Alignment and Equilibrium
Given a vector-labeled graph $(G,\{\boldsymbol{v}_{i}\}_{i\in V})$ associated with the state $\boldsymbol{\theta}$, we reformulate the equilibrium condition (3) (equivalently (4)) in geometric terms.
@@
where $\boldsymbol{v}_{i}:=(\cos\theta_{i},\sin\theta_{i})$.
@@
The condition (5) is equivalently written as
$\sum_{j\in N(i)}e^{\mathrm{i}\theta_{j}}=\mu_{i}e^{\mathrm{i}\theta_{i}}.$ (6)
Indeed, rewrite all vectors $\boldsymbol{v}_{i}$ in (5) in terms of their Cartesian components, yielding
@@
The condition (7) is equivalent to the equality between the following two complex numbers:
@@
which can be compactly rewritten as (6) using the Euler formula $e^{I\theta_{i}}=\cos\theta_{i}+I\sin\theta_{i}$.
Now that we have derived the equivalence between conditions (5) and (6), our goal reduces to proving the equivalence between (6) and (3), namely
$\forall i\in V,\qquad\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}}\Longleftrightarrow\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})=0.$
The implication from the left-hand side to the right-hand side is established by showing that
@@
This can be verified by a direct computation:
$I\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})+\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})$
@@
To prove the other direction, we construct for every node $i\in V$, such a $\mu_{i}$ that satisfies the left-hand side. Dividing both sides of (6) by $e^{I\theta_{i}}$ (noting that $e^{I\theta_{i}}\neq 0$), we obtain
$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}=\mu_{i}.$
Compute
$\frac{\sum_{j\in N(i)}e^{I\theta_{j}}}{e^{I\theta_{i}}}$ $=\sum_{j\in N(i)}e^{I(\theta_{j}-\theta_{i})}$
$=\sum_{j\in N(i)}(\cos(\theta_{j}-\theta_{i})+I\sin(\theta_{j}-\theta_{i}))$
$=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})$
where the last equality is due to the fact $\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})=0$. Thus
$\sum_{j\in N(i)}e^{I\theta_{j}}=\mu_{i}e^{I\theta_{i}}$
holds for every $i\in V$ by letting $\mu_{i}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})$. ∎
@@
4.3. Angles and Second-Order Stationarity
Given an equilibrium $\boldsymbol{\theta}\in\mathbb{R}^{n}$ of the Kuramoto model on graph $G=(V,E)$, if there exists a node $i\in V$ whose corresponding $\mu_{i}<0$, then $\boldsymbol{\theta}$ cannot be a second-order stationary point of the Kuramoto energy (2). This is because a small perturbation that moves $\boldsymbol{v}_{i}$ toward the direction of $\sum_{j\in N(i)}\boldsymbol{v}_{j}$ strictly decreases the energy.
Lemma 4.4.
If $\boldsymbol{\theta}\in\mathbb{R}^{n}$ is a second-order stationary point of (2), then for every $i\in V$ there exists $\mu_{i}\geq 0$ such that
$\sum_{j\in N(i)}\boldsymbol{v}_{j}=\mu_{i}\boldsymbol{v}_{i}.$
In particular, either $\sum_{j\in N(i)}\boldsymbol{v}_{j}=\boldsymbol{0}$ (i.e., $\mu_{i}=0$), or $\angle\!\left(\sum_{j\in N(i)}\boldsymbol{v}_{j},\,\boldsymbol{v}_{i}\right)=0$.
Proof of Lemma 4.4.
By Proposition 3.4, the diagonal entry of the Hessian satisfies
$\boldsymbol{H}_{ii}=\sum_{j\in N(i)}\cos(\theta_{i}-\theta_{j}).$
Since $\theta$ is an equilibrium, we have
$\sum_{j\in N(i)}\sin(\theta_{j}-\theta_{i})=0,$
which implies that
$\sum_{j\in N(i)}e^{\mathrm{i}(\theta_{j}-\theta_{i})}=\sum_{j\in N(i)}\cos(\theta_{j}-\theta_{i})\in\mathbb{R}.$
By Lemma 4.1, this real quantity equals $\mu_{i}$ at first-order stationary point $\boldsymbol{\theta}$. Hence $\boldsymbol{H}_{ii}=\mu_{i}$. If $\mu_{i}<0$, then $\boldsymbol{H}_{ii}<0$, contradicting the positive semidefiniteness of $\boldsymbol{H}$. ∎
@@ 4.3. Angles and Second-Order Stationarity
Figure 6. The solid arrow represents $\boldsymbol{v}_{i}$, and the dashed arrow represents $\sum_{j\in N(i)}\boldsymbol{v}_{j}$. For each node $i$, the four panels show all possible relations between $\boldsymbol{v}_{i}$ and $\sum_{j\in N(i)}\boldsymbol{v}_{j}$. Among them, case (A) violates the first-order condition, while case (B) violates the second-order stationary condition, since the aggregated phasor $\sum_{j\in N(i)}\boldsymbol{v}_{j}$ lies outside the feasible (gray) region corresponding to $\mu_{i}\geq 0$.
@@ 5.1. A Fundamental Geometric Fact
Let $\boldsymbol{v}_{a},\boldsymbol{v}_{b}\in\mathbb{S}^{1}$. Suppose there exist a vector $\boldsymbol{q}\in\mathbb{R}^{2}$ and scalars $\mu_{a},\mu_{b}\in\mathbb{R}$ such that
@@
Case 1: $\mu_{a}\neq-1$ and $\mu_{b}\neq-1$. In this case, from (9), it must hold that
@@
This leads to two subcases:
(1.1) If $\mu_{a}=\mu_{b}\neq-1$, then $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$;
(1.2) If $\mu_{a}+1=-(\mu_{b}+1)$, then equivalently, we have $\mu_{a}+\mu_{b}=-2$. This implies $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$.
Case 2: $\mu_{a}=\mu_{b}=-1$. In this case, we have
$\boldsymbol{q}+\boldsymbol{v}_{a}+\boldsymbol{v}_{b}=\boldsymbol{0}.$
@@
Figure 7. The feasible region for $(\mu_{a},\mu_{b})$ consists of: (1) the black line $\mu_{a}=\mu_{b}$, excluding the intersection point, corresponding to $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$; (2) the red line $\mu_{a}+\mu_{b}=-2$, excluding the intersection point, corresponding to the antipodal case $\boldsymbol{v}_{a}=-\boldsymbol{v}_{b}$; (3) their intersection point (1,1)(-1,-1), marked by a hollow dot, at which no additional constraint is imposed on the vectors $\boldsymbol{v}_{a}$ and $\boldsymbol{v}_{b}$.
@@
With $\mu_{a},\mu_{b}\geq 0$, Lemma 5.1 leaves only case (1.1), hence $\boldsymbol{v}_{a}=\boldsymbol{v}_{b}$. ∎
We introduce a geometric analogue of the notion of twins in graph theory, tailored to the Kuramoto model.
Definition 5.3 (Geometrically closed twins).
Motivated by the notion of closed twins in graph theory, we say that, at a state $\boldsymbol{\theta}$ of the Kuramoto model on $G$, nodes $i$ and $j$ are called geometrically closed twins if their phasors $\boldsymbol{v}_{i}$ and $\boldsymbol{v}_{j}$ satisfy the relations of Lemma 5.1.
Definition 5.4 (Geometrically stable closed twins).
We say that, at a state $\boldsymbol{\theta}$ of the Kuramoto model on $G$, nodes $i$ and $j$ are called geometrically stable closed twins if their phasors $\boldsymbol{v}_{i}$ and $\boldsymbol{v}_{j}$ satisfy the relations of Lemma 5.1 and, in addition, $\mu_{i},\mu_{j}\geq 0$.